Need help with this, The inlet pipe of an oil tank can fill the tank in 1 hour, 30 minutes. The outlet pipe can empty the tank in 1 hour. How long does it take to empty the tank if both pipes are open?
Answer by venugopalramana(3286)

The inlet pipe of an oil tank can fill the tank in 1 hour, 30 minutes. The outlet pipe can empty the tank in 1 hour. How long does it take to empty the tank if both pipes are open?
inlet pipe fills 1 tank in 1hr 30 mts =1.5=3/2 hrs.
or inlet pipe fills ...1/(3/2)=2/3 tank in 1 hr.
out let pipe empties 1 tank in 1 hr.
...so assuming the tank is full at the start water to be emptied =1 tank
and in 1 hr inlet fills 2/3 tank and out let empties 1 tank ..so net out flow is 1-2/3=(3-2)/3=1/3 tank in 1 hour.
so 1/3 tank emptying needs .....1 hr
then 1 tank emptying needs ......? hrs...= 1*1/(1/3)=3 hrs.